OpenStudy (anonymous):

Write a quadratic equation that fits the points (0, -5), (1, 3), and (5, -5).

OpenStudy (anonymous):

do you know what the quadratic equation is?

OpenStudy (anonymous):

Mhhmm

OpenStudy (anonymous):

plug in for abc

OpenStudy (anonymous):

This question is multiple choice: f(x) = -2x^2 + 10x - 5 f(x) = -1.5x^2 + 9.5x - 5 f(x) = (1/2)x^2 - (-5/2)x - 5 f(x) = 8x - 5

OpenStudy (anonymous):

So ax^2 + bx + c = 0

OpenStudy (anonymous):

@mocham

OpenStudy (anonymous):

Oh, haha:) it's a

OpenStudy (anonymous):

=)